The following table summarizes the amount of time customers had to wait in line at a bank:

Wait Time # customers
1 min 7
2 min 8
3 min 6
4 min 3
5 min 2
6 or more min 9

If a customer is selected at random, what are the probabilities that:

a) The wait time is over 3 minutes

b) The wait time is exactly 2 minutes

c) The wait time is at least 2 minutes
Answer #1

Total number of customers = 35

a) Number of customers with wait time over 3 minutes = 3+2+9 = 14

Probability = 14/35 = 0.4

b) Number of customers with wait time exactly 2 minutes = 8

Probability = 8/35 = 0.22857

c) Number of customers with wait time at least 2 mins = No. of customers with wait time of 2 or more than 2 mins = 8+6+3+2+9 = 28

Probability = 28/35 = 0.8